Skybet 1350 Cheltenham money back if horse loses

  • 8 runners, 1 very clear fav. Very generous.

  • New to matched betting?

    My Matched Betting Academy is the best place to get started. Learning the fundamentals takes 10-15 minutes and you’ll make £15 in the process.

    Learning the fundamentals takes 10-15 minutes and you’ll make £15 in the process.

    LJH 2

    THANK YOU Skybet. Needed that after a horrid week doing horse promos. 15 races and hit only 1 £25 refund.

    +0
    Dbb 4

    love skybet :)

    Since it is a money back on loss offer I have put my back in tonight and I will see what lay prices are like tomorrow and hope for some drops.

    worst case it goes the other way and I make a couple of quid less guaranteed profit, cant actually lose on this one

    +0
    Tony 18

    Love this offer :D

    Overlay city with that favourite!

    +0
    beh 16

    Also be sure to opt in to their racing rebate, 20% of losses back (10x WR) – https://www.skycasino.com/promotion/38269/racing-rebate

    +0
    Pavlon 3

    Yeah, they actually emailed me about this as market research a couple of weeks ago. Seems they’re going to run this offer for the first race each day at Cheltenham.

    +0
    greig A 2

    I’m thinking of having a bit of a punt on one of the 2nd or 3rd favs, if it doesn’t come in i’m only a couple of quid down, or maybe just lay underlay it? Anyone else see this as a bit of a risk free gamble?

    +0
    Pavlon 3

    (actually, that might have been for the Cheltenham Festival in mid-March)

    Greig – yeah, think I’ll underlay too. Just lay £15.

    Good luck

    +0
    FoG_BLoG 61

    Im going €25 on Smad Place at 7/1.

    I figure 10% chance I will have €200 and 90% chance a free bet of €25 which is worth €20 in real money.

    Better than having 55% chance of returning €39.29 and only 45% chance at the free bet.

    Besides if Will Hill have their money back second tomorrow too I will probably have the favourite covered off in that one.

    +0
    beh 16

    How do you figure? Seems you’ve plucked your percentages from nowhere.

    If you want to have a punt, fair enough.

    +0
    Ice Magic 5

    Well I’ve gone Smad place. Guaranteed £14
    Thanks very much sky

    +0
    FoG_BLoG 61

    Sorry, I didn’t pluck the figures from nowhere. I convert the bookies odds to implied probabilities. So 4/7 becomes 7/(4+7) = 0.64 and 7/2 becomes 1/(7+1) = 0.125. Obviously the bookie implied probabilities are not accurate because they build in a bias. So I divide by 1.2 to get an adjusted implied probability.

    Its not fair on the favourite there though, less of an edge on the short odds, so he is probably close to around 60% chance of winning.

    I have a lot of respect for the bookie’s and their odds. They have banks of stats and tons of experience and make a shocking amount of money – I know alot of ppl don’t consider their odds as an accurate view of the actual probability. I consider them the masters of risk and probabilities.

    +0
    MarkCorrigan 17

    Haha! Masters of risk and probabilities. Why aren’t you just using the exchange odds as a guide for implied probabilities? (the market is perfectly competitive and doesn’t suffer from favourite-longshot bias etc.) The favourite is 1.69 currently – 59% chance of winning.

    +0
    FoG_BLoG 61

    Its not really important how you estimate the probs. It looks like the best punt is on one of the longer shots anyway. That’s all I’m saying – if you want to punt it.

    But strictly matched betting go with the one which gives the tightest odds between exchange and bookie.

    Just for me I want to punt this especially as I’ll have the favourite backed on will hill for their offer.

    +0
    kjillman 1

    same here ice magic, Smad place -of course placed £2 with another bookie in case it wins

    +0
    MarkCorrigan 17

    You’re actually using the offer optimally… don’t worry, FoG. Assume 80% free bet extraction.

    If you back the favourite at 1.57 (and using the exchange odds of 1.69 to calculate implied probability), your expected gain is roughly £6.39.

    If you back Smad Place at 8 (and using exchange odds of 9.6 for subjective probability), your expected gain is roughly £13.75. You can of course choose to lay this if you wanted and lock in a little bit less than £13.75 after commission.

    +0
Viewing 15 replies - 1 through 15 (of 44 total)

You must be logged in to reply to this topic.